# Oversmoothing, “Oversquashing”, Heterophily, Long-Range, and more: Demystifying Common Beliefs in Graph Machine Learning

### v1

@@ 2 Is Oversmoothing Really a Problem?
We show a simple training scenario where we see how OSM is not a property of all DGNs and how different elements make it difficult to draw clear conclusions. We train several DGNs under two propagation variants: the vanilla $AXW$ update and the rescaled $AX(2W)$, inspired by Roth and Liebig [81, Figure 1]. We measure OSM with 2 different metrics: Dirichlet Energy (DE) and its norm-normalized version, the Rayleigh Coefficient (RQ), which was also previously used in some works [75, 39, 81, 68]. Figure 1 shows three key facts: (i) some architectures never collapse, (ii) a minor rescaling can reverse the trend, and (iii) DE and RQ often disagree. Hence OSM is neither universal nor straightforward to diagnose.
@@
Figure 1: (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}=\mathrm{tr}(X^{T}\Delta X)$ and the feature norm $\|X\|_{F}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}=\mathrm{tr}(X^{T}\Delta X)/\|X\|$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds. A larger version for better visualization is available in Fig. 8.
@@
Regarding the first statement, although it is true that if there is total collapse to the same value then the embeddings will not be informative at all, the main problem remains the node embeddings’ separability. As shown in the previous section, some changes in the architectures can avoid OSM, but they might have no impact on the overall accuracy. For instance, multiplying by two the weight matrix leads to a general increase in DE for all architectures, however, the accuracy will remain the same since the embeddings have been simply scaled up and the embeddings’ separability is not affected negatively. In addition, avoiding an embedding collapse does not necessarily lead to an improvement in accuracy. For instance, comparing a GCN with and without bias, both versions show a decrease in performance as the number of layers increases, whereas only GCN shows a collapse in DE [83, Figure  3].
@@
In conclusion, low accuracy in DGNs cannot be attributed to OSM alone. The separability of node embeddings plays a major role, where other training problems such as vanishing gradients or over-fitting also arise when using a big number of layers [107, 27, 103, 7].
@@ 4.1 Belief: Oversquashing as Synonym of Topological Bottleneck
Another, slightly more technical way to see why low Jacobian sensitivity does not imply the presence of any topological bottlenecks is to follow the chain of upper bounds that link the metrics used to measure the computational and topological bottlenecks [16, 39, 52, 6]. Several recent works [16, 29] have shown that sensitivity is upper bounded by a term that includes the effective resistance, a purely topological distance metric that quantifies the expected commute time of a random walk between nodes $u$ and $v$ [55]. In particular, the bound subtracts the lower bound on the maximum effective resistance, which depends on the inverse of the spectral gap (a proxy for topological bottlenecks) [21, 63]. This connection provides a useful intuition: as the topological bottleneck gets worse, the lower bound on the maximum effective resistance increases, which in turn reduces the upper bounds on the sensitivity of Black et al. [16]. However, the converse does not hold. There are graphs where the maximum effective resistance between distant nodes is large despite the absence of topological bottlenecks. For instance, consider the example of a grid graph in Figure 4 (left) where there is no topological bottleneck, yet the effective resistance between diagonally opposite corners grows linearly with the grid size. As a result, sensitivity between those nodes decays with depth, even though the graph has no identifiable topological bottlenecks. This illustrates that computational bottlenecks can arise independently of topological ones, and that low sensitivity does not necessarily imply the presence of either of them.
@@ Appendix B Sensitivity Decreases on a Grid Graph without Topological Bottlenecks
Appendix C OSM does not always happen

### v2

@@ 2 Is Oversmoothing Really a Problem?
We show a simple training scenario where we see how OSM is not a property of all DGNs and how different elements make it difficult to draw clear conclusions. We train several DGNs under two propagation variants: the vanilla $AXW$ update and the rescaled $AX(2W)$, inspired by Roth and Liebig [82, Figure 1]. We measure OSM with 2 different metrics: Dirichlet Energy (DE) and its norm-normalized version, the Rayleigh Coefficient (RQ), which was also previously used in some works [18, 39, 82, 68]. Figure 1 shows three key facts: (i) some architectures never collapse, (ii) a minor rescaling can reverse the trend, and (iii) DE and RQ often disagree. Hence OSM is neither universal nor straightforward to diagnose.
@@
Figure 1: (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}=\mathrm{tr}(X^{T}\Delta X)$ and the feature norm $\|X\|^{2}_{F}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}=\mathrm{tr}(X^{T}\Delta X)/\|X\|^{2}_{F}$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds. A larger version for better visualization is available in Fig. 8.
@@
Regarding the first statement, although it is true that if there is total collapse to the same value then the embeddings will not be informative at all, the main problem remains the node embeddings’ separability. As shown in the previous section, some changes in the architectures can avoid OSM, but they might have no impact on the overall accuracy. For instance, multiplying by two the weight matrix leads to a general increase in DE for all architectures, however, the accuracy will remain the same since the embeddings have been simply scaled up and the embeddings’ separability is not affected negatively. In addition, avoiding an embedding collapse does not necessarily lead to an improvement in generalization accuracy. For instance, comparing a GCN with and without bias, both versions show a decrease in performance as the number of layers increases, whereas only GCN shows a collapse in DE [84, Figure  3].
@@
In conclusion, low accuracy in DGNs cannot be attributed to OSM alone. The separability of node embeddings plays a major role, where other training problems such as vanishing gradients or over-fitting also arise when using a big number of layers [108, 27, 104, 7, 76].
@@ 4.1 Belief: Oversquashing as Synonym of Topological Bottleneck
Another, slightly more technical way to see why low Jacobian sensitivity does not imply the presence of any topological bottlenecks is to follow the chain of upper bounds that link the metrics used to measure the computational and topological bottlenecks [16, 39, 52, 6]. Several recent works [16, 29] have shown that sensitivity is bounded by above by a term that includes the effective resistance, a purely topological distance metric that quantifies the expected commute time of a random walk between nodes $u$ and $v$ [55]. In particular, the bound subtracts the lower bound on the maximum effective resistance, which depends on the inverse of the spectral gap (a proxy for topological bottlenecks) [21, 63]. This connection provides a useful intuition: as the topological bottleneck gets worse, the lower bound on the maximum effective resistance increases, which in turn reduces the upper bounds on the sensitivity of Black et al. [16]. However, the converse does not hold. There are graphs where the maximum effective resistance between distant nodes is large despite the absence of topological bottlenecks. For instance, consider the example of a grid graph in Figure 4 (left) where there is no topological bottleneck, yet the effective resistance between diagonally opposite corners grows linearly with the grid size. As a result, sensitivity between those nodes decays with depth, even though the graph has no identifiable topological bottlenecks. This illustrates that computational bottlenecks can arise independently of topological ones, and that low sensitivity does not necessarily imply the presence of either of them.
@@ Appendix B Sensitivity Decreases on a Grid Graph without Topological Bottlenecks
Appendix C Oversmoothing Does Not Always Happen